### v2

# Beyond prompt brittleness: Evaluating the reliability and consistency of political worldviews in LLMs

Abstract
Due to the widespread use of large language models (LLMs) in ubiquitous systems, we need to understand whether they embed a specific “worldview” and what these views reflect. Recent studies report that, prompted with political questionnaires, LLMs show left-liberal leanings Feng et al. (2023); Motoki et al. (2024). However, it is as yet unclear whether these leanings are reliable (robust to prompt variations) and whether the leaning is consistent across policies and political leaning. We propose a series of tests which assess the reliability and consistency of LLMs’ stances on political statements based on a dataset of voting-advice questionnaires collected from seven EU countries and annotated for policy domains. We study LLMs ranging in size from 7B to 70B parameters and find that their reliability increases with parameter count. Larger models show overall stronger alignment with left-leaning parties but differ among policy programs: They evince a (left-wing) positive stance towards environment protection, social welfare state and liberal society but also (right-wing) law and order, with no consistent preferences in foreign policy and migration.
1 Introduction
Ensuring space for a plurality of ideas and opinions is crucial to keep democracies healthily operating (McQuail, 1993; Loecherbach et al., 2020). Given that one such space — the interaction with chatbots powered by LLMs — is becoming increasingly prevalent, the importance of unveiling biases in LLMs’ output is clear. In this context, people can easily ask chatbots to perform a variety of tasks such as generating arguments, summarizing and writing an email. Investigating how heterogeneous the worldviews embedded in these LLMs is, therefore, of extreme relevance for several downstream tasks, not only chatbot interactions. Continuing previous work on identifying biases in general NLP resources and models (Hovy and Prabhumoye, 2021), a number of studies have found biases of numerous types in LLMs, relating to gender Kotek et al. (2023), race (Omiye et al., 2023), culture Arora et al. (2023); Wang et al. (2023b), and political position Feng et al. (2023). Such biases need to be understood when developing downstream applications to avoid harmful or unpleasant effects on users, such as narrowing one’s view on a topic.
In this paper, we focus on a specific type of bias, namely political bias in LLMs. Recent studies found that LLMs show a bias in that they agree more with left-wing political positions (Feng et al., 2023; Motoki et al., 2024). However, the scope and interpretation of these findings is not clear yet: Political positioning is an inherently multidimensional phenomenon, and while political individuals and organizations (e.g., parties) typically exhibit substantial (even if typically imperfect) internal consistency (Moskowitz and Jenkins, 2004; Tavits, 2007), this is not necessarily true for LLMs, which have only a weak notion of consistency Basmov et al. (2024). This opens two possible interpretations (and a sliding scale between them) for political bias in LLMs, which it would arguably be instructive to distinguish: (a), LLM show preferences primarily at the level of individual statements without major consideration to other, related statements; this would amount to stance taking (Küçük and Can, 2020) with respect to individual targets. Alternatively, the studies can be taken to indicate that (b), LLMs exhibit political belief systems with a substantial amount of internal consistency, similar to human political individuals – in a word, a political “worldview”, a term which is used for this concept in political science Ecker et al. (2021) and has also been suggested to apply to LLMs (Bender et al., 2021).
We suggest that interpretations (a) and (b) can be distinguished with the help of two diagnostics that we take to be constituted of “worldviews”. First, the models should show high reliability in assessing political statements [^1] We adopt the term “reliability”, as consistency over testing replications, from psychometry (American Educational Research Association et al., 1999). — that is, they should give consistent answers irrespective of the formulation of the prompts. If this is not the case, models merely react to linguistic peculiarities, namely lexical choice, token or position biases (see Section 2 for details). Second, the models should show consistency in their political position, that is, should exhibit a consistent stance towards broad policy domains, with limited variance among statements within these domains, as well as a consistent commitment to a right or left leaning across domains.
To improve our understanding of political bias in current LLMs, we make three contributions:
  1. 1.
    We release ProbVAA, a dataset with statements on policy measures from seven EU countries with the answers (stance) from several political parties annotated on policy domains. We expand ProbVAA with paraphrased, negated, and semantically inverted versions of the core statements (§4).
  2. 2.
    We propose a method for evaluating the reliability of the LLMs’ output across variations of statements and prompts (§3). It adheres to psychometric standards and involves expanding the dataset in accordance with these principles. This work is most similar to Shu et al. (2023), but prioritizes a data-centric approach, indicating that the analysis can be conducted on both open- and closed-source models, solely utilizing the responses produced by the LLM.
  3. 3.
    We evaluate a range of SOTA LLMs on the ProbVAA dataset, finding substantial differences among LLMs with regard to reliability (§6). When evaluating stance on reliable statements (§7), we find that LLMs align more with left-leaning parties overall, but lack consistency regarding leanings: they tend to have no preference for some domains (migration, foreign policy) but agree with policies as divergent as pro-environment and law and order.
2 Related Work
Political Positioning.
The characterization of political positions is an important topic in political science, and a considerable number of computational models has shown that positions can be inferred from political texts (e.g., Laver et al., 2003; Slapin and Proksch, 2008; Glavaš et al., 2017). Comparing the positioning of political parties at low dimensional level under pre-defined scales remains an elusive goal in political science Heywood (2021). One of the most widely used scales is left-right, arguably distinguishing between progressive position (left), conservative positions (right) and compromise positions (center). Despite concerns about its validity (Kitschelt, 1994; Jahn, 2023), the scale has been validated broadly across countries (Evans et al., 1996; Budge et al., 2001) and also formed the basis for previous analyses of political bias in LLMs. An alternative to positioning actors on a global scale is to carry out a fine-grained analysis at the level of individual policy domains (Iversen, 1994; Ceron et al., 2023). For our consistency analysis in Section 7, we look at both of these levels (left-right scale and positioning within policy domains).
Worldviews in LLMs.
Recent work has examined LLMs’ political ideology using global surveys such as Political Compass (Feng et al., 2023; Motoki et al., 2024; Rutinowski et al., 2023), or more country-specific questionnaires such as Pew Research’s ATP, World Values Survey (Durmus et al., 2022), and voting advice applications (VAAs) (Hartmann et al., 2023).
Different methods have been utilized to capture bias, including integrating the agreement options directly within the prompt, averaging model responses Rutinowski et al. (2023) and prompt paraphrases Feng et al. (2023). Another approach stream leveraged the form of multiple-choice questions where the response polarity was determined by extracting log-probabilities of answer options to obtain the model’s opinion distribution Santurkar et al. (2023), shuffling the option order within the prompt Durmus et al. (2023) and using response sampling with randomizing question order Motoki et al. (2024). However, each approach tackled a single aspect of reliability – either the LLM’s prompt sensitivity or the stability of their output.
LLM Probing.
The assessment of output variability and the quantification of model reliability in recent studies have involved the application of psychometric methods from social psychology. These studies have utilized standardized methodologies Dayanik et al. (2022) and questionnaires to create controlled environments for extracting reliable “attitudes” from LLMs (Tjuatja et al., 2023; Dominguez-Olmedo et al., 2023; Shu et al., 2023). Such approaches have proven to be instrumental in examining various societal biases in LLMs (Arora et al., 2023; Wang et al., 2023b; Hada et al., 2023; Esiobu et al., 2023; Shu et al., 2023). However, the exploration of psychometric methods to investigate political bias remains limited.
LLM Brittleness.
There is a growing awareness of the fact that the input to an LLM is not the only, and maybe not even the most important, factor determining the output. For example, Min et al. (2022) show that swapping out gold labels for random ones only slightly reduces performance – a pattern that remains stable across almost all tested models regardless of the prompt template used. Khashabi et al. (2022) observe that soft prompts manage to solve a task even when presented as an arbitrary text, staying surprisingly close (within a 2% range) to the best soft prompt of the same size designed for that specific task. Finally, the meaning of prompts can be overshadowed by the choice of target words Webson and Pavlick (2022) which goes hand-in-hand with observed high result variance caused by recency and common token bias phenomena when the model chooses the most frequent token Zhao et al. (2021), or position bias when the model prioritizes labels that appear at a specific position Zheng et al. (2023).
3 Reliability-Aware Bias Analysis
Figure 1: The workflow for creating model inputs. The procedure for augmenting original statements is described in § 4.1, and prompt design is described in § 5.2.
Following up on this motivation, we now present our framework for evaluating the political bias of LLMs which involves two key elements: (1) enrichment of the dataset with prompt variations and policy-domain annotations and (2) evaluation of the reliability of answers and the consistency of stances.
Figure 1 illustrates the workflow for creating model inputs. Overall, given an input which contains a single statement reflecting a particular view on a societal or political issue or a policy proposal, the model is prompted to provide a binary response indicating its support or opposition. In the subsequent discussion, we refer to model response as binarized free-text response with agreement/approval as opposed to disagreement/disapproval towards the given input.
After collecting our target dataset (details in § 4.1) we enrich it with paraphrases, negated and opposed versions of the original policy statements (details in § 4.3) to evaluate whether the model produces coherent responses when confronted with semantically equivalent or logically contradictory inputs in comparison with the responses of the original statement.
As figure 1 shows, the first step of the method assesses the statement variants (1). In addition to that, the reliability with respect to variations of prompt templates is evaluated by (2) using two types of instructions (personal and impersonal questions), (3) using synonyms for the response alternatives that the model should select, and (4) swapping the order of the alternatives (§ 5.2).
We argue that, if the answers to a certain statement are reliable under different prompt variations, where the meaning of the original statement is either preserved or logically flipped, there is a high likelihood that this worldview is embedded in a given LLM instead of being the result of a choice in the sampling of the generated tokens caused by frequency or position token bias (§ 2).
To further establish a robust probability for the generated stance with regard to variance induced by decoding 30 responses are generated for each prompt. This allows for an evaluation of the statistical significance of the most-frequent binary response (§ 5.4).
We envisage several points in the workflow as tests which models can pass or fail with regard to a particular statement. As illustrated in figure 2, the test types are (1) robustness to sampling (with a fixed prompt), (2) robustness to paraphrasing/negation/semantic inversion of the original statement, and (3) robustness to label-order inversion in the prompt template. Only statements on which the models pass all tests are used to assess the models’ attitudes. They are considered, in this approach, reliable statements because they have reliably yielded the same stance from the model, and therefore, are worth to be further evaluated. Policy-domain annotations on the dataset make it possible to make the analysis of the reliable statements more fine-grained (§ 6 and 7).  [^2] We make the augmented dataset, including all tests, the models’ responses and code, available here https://github.com/tceron/eval_political_worldviews.
Figure 2: Overview of reliability tests
4 The ProbVAA Dataset
4.1 Sources
To assess the potential political worldviews embedded in LLMs, we collect a set of statements derived from Voting Advice Applications (VAAs). VAAs are tools that provide voters with insights on which parties are best aligned with their own opinions regarding policy issues. Unlike the frequently used Political Compass questionnaire, which categorizes political attitudes into a two-axis system (left/right and authoritarian/libertarian), VAAs offer a nuanced approach that ground political leanings in stances towards practical policies Palfrey and Poole (1987); Tavits (2007). These stances allow for a direct comparison of responses with those from national parties and/or candidates. On the one hand, this offers a more unbiased basis for measuring political leanings, as it does not rely on the questionnaire designer’s external classification to determine if an answer aligns with the "left" or "right" side of the political spectrum. On the other hand, it covers a wide range of policy domains that varies from environmental protection to government expenditures, providing more fine-grained insights on the types of biases.
Concretely, we collect the statements and answers of VAAs of the parliamentary elections (ranging from 2021 to 2023) from seven countries (Poland, Hungary, Italy, Germany, Netherlands, Spain, and Switzerland) in 7 languages. The length of questionnaires varies between 20 and 60 questions (a breakdown of the number of statements per country is shown in Figure 8, Appendix C).
Most of them are in the format of statements, except for the Swiss VAA, which contains questions that we manually convert to statements to align with the other countries. The dataset contains a total of 239 unique statements in the source languages (Switzerland has 60 statements for each language – German, Italian, and French – but only 60 count as unique given that they are the same statements). In order to answer our research questions, we annotate the datasets in a number of ways discussed below.
4.2 Policy-Domain Annotation
Category $\kappa$
Open foreign policy 0.85
Liberal economic policy 0.78
Restrictive financial policy 0.65
Law and order 0.58
Restrictive migration policy 0.88
Exp. environment protection 0.79
Exp. social welfare state 0.72
Liberal society 0.73
Table 1: Fleiss $\kappa$ between three annotators for policy-domain annotations.
We have enriched ProbVAA with policy-domain annotations based on the pattern of the Swiss VAA, SmartVote [^3] More info on https://www.smartvote.ch/en/wiki/methodology-smartspider/23_ch_nr?locale=en_CH. It contains annotations that allow for the visualization and deeper understanding of the positioning of parties according to predominant policy-domains in the political spectrum. We draw from the documentation provided by SmartVote where eight categories (considered stances on policy domains) are defined: open foreign policy, liberal economic policy, restrictive financial policy, law and order, restrictive migration policy, expanded environmental protection, expanded social welfare state, and liberal society. These categories are based on domains identified in Swiss political spectrum Hermann and Leuthold (2001, 2003), but that are generalized across European countries, as evidenced by the similarity with domains analyzed in cross-European studies such as the Chapel Hill Survey Jolly et al. (2022).
When answering ‘agree’ to a statement emphasizes any of the eight given policies, the statement is marked as a ‘agree’ with that policy domain, while disagreements with a policy are annotated as ‘disagree’. Three annotators with background in traditional or computational political science extended the annotations to the other countries. Table 1 shows that inter-annotator agreement – which is calculated with agreement between ‘agree’, ‘agree’ and no label per statement – is good. The final gold annotations are drawn from the majority votes. Note that some statements do not fall into any category. Therefore, the gold annotations contain 193 statements in total (tables 9 and 10, Appendix A provide examples and details).
4.3 Robustness to Statement Variations
We introduce three variants of each policy statement to test the models’ reliability (cf. statement variants, Figure 1 and robustness to statement variations in Figure 2).
Reliability under paraphrasing
With paraphrasing, we aim to measure how consistently the models (or humans) generate the same stance on semantically similar statements. For every statement ($S$) in the source language ($S_{src}$) and in English ($S_{en}$), we generated three paraphrases using ChatGPT4. Native speakers read a sample of 60 paraphrases for 20 $S$s in the source language and confirmed that they are syntactically and semantically correct.
Reliability under negation and semantic opposite
These two tests evaluate whether the models (or humans) generate the opposite stance when presented with a negated or semantically inverted version of the original policy statement, i.e., agree for the original and disagree for the opposite and vice versa). Given statement $S$, its negated opposite, which we denote as $\text{Neg}(S)$ is its logical opposite, which is constructed by adding an overt negation marker in the appropriate position in the statement.
The other type, which we call semantic opposite and denote $\text{Opp}(S)$, is a statement that takes the semantically opposite sense to the original one while not using an overt negation marker. A minimal number of words is modified to convert the semantic meaning of the sentence.
Each statement in the source language is annotated by a native speaker. Annotators are asked to create $\text{Neg}(S)$ by adding a marker corresponding to ‘not’ or ‘don’t’ in the source languages. As for $\text{Opp}(S)$, annotators are instructed to first try modifying the head verb in the statement or, if this is not possible, the focal adjective. If neither can be altered, they are asked to apply the minimal change necessary to invert the sentence’s meaning.
Translations
Every statement ($S$) with their respective $\text{Neg}(S)$ and $\text{Opp}(S)$ has been automatically translated into English with the commercial translation tool DeepL. The quality of the translations has been validated on a subset of the statements by the authors. Altogether, this results in 1434 statements in English and in the source languages. The ProbVAA dataset consists of both English and original-language statements, but we only use the translated statements for this study because the evaluated models have not been instruction fine-tuned in the source languages.
5 Experimental Setup
In this section, we describe the models that we use (§ 5.1), our prompting, sampling, and output-mapping strategy (§§ 5.25.4).
5.1 Models
Given that we formulate our prompts as zero-shot instruction, we opt for the instruction-following model families that vary in parameter size and contain different sets of pretraining data  [^4] We also evaluated the base models, but they do not produce any reliable answers, and are therefore not reported here.. Specifically, we focus on all size variants of LLama-2-Chat (7B, 13B, 70B) Touvron et al. (2023), the XXL variant of Flan-T5 (11B, Chung et al., 2022), Mistral Instruct (7B, Jiang et al., 2023) for open-sourced models and GPT-3.5 as a closed-source model. The models form three natural classes: small ($<$10B parameters, mistral-7b and llama2-7b), medium (between 10 and 19B (flanT5-xxl-11b and llama2-13b) and big ($>$20B, gpt3.5-20b and llama2-70b). All models utilize the top-p nucleus decoding approach, suggested by Holtzman et al. (2020), when generating responses, aiming to capture the model’s stance distribution through the sampling of the output (Cf. Appendix B for information on the implementation).
5.2 Prompt Design
For clarity, we define a prompt as composed of two parts, the prompt template (which contains the instruction given to the model) and the statement (an instance from ProbVAA).
When designing model inputs, we aim at creating templates that reliably elicit responses from models that can be clearly aligned with a defined stance, so either positive or negative. [^5] An example of an invalid response is I don’t know or I don’t have personal opinions. Considering recent research findings indicating that the meaning of prompts can be overshadowed by the choice of the target words as discussed in § 2, we aim at diversifying the choice for prompt templates from various angles. We distinguish between personal and impersonal templates (cf. prompt variants as shown in Figure 1). Personal templates formulate the instruction as a direct question (e.g. Do you agree or disagree?, Are you pro or con?), whereas impersonal are framed as objective tasks (Analyze the following statement into the labels “favorable” or “detrimental”, Classify the following statement as…). The context for evaluating the prompts is specified as Consider the long-term societal impact…. Additionally, we vary the wording of the stance (e.g. favorable, detrimental, advantageous, disadvantageous, support, oppose) to explore potential model biases in responding to specific wordings (cf. semantic label order, Figure 1). After a pilot experiment to test which prompts elicit most valid responses, we selected 3 personal and 3 impersonal prompt templates among 8 impersonal and 6 personal templates (Appendix B.1 details the selection process). Refer to the implemented prompt templates in Table 7, Appendix A.
Reliability under inverted labels
In order to test sensitivity of the models to subtle template changes each template is furthermore presented in two versions: the original one and the version where the order of the labels is swapped, e.g., if a template states, Analyze the following statement into the labels “favorable or detrimental”, the inverted-label version corresponds to “detrimental or favorable”. A reliable model is expected to yield the same response independent of label order (cf. robustness to label-order inversion, Figure 2).
Reliability under varied templates
In addition to altering the statements, we modify the templates to investigate if the model maintains consistent stances with semantically equivalent templates. Previous research has demonstrated the impact of template variation on the results (Min et al., 2022; Khashabi et al., 2022). We hypothesize that variations in templates are likely to be an influential factor in shifts in the models’ generated stance.
5.3 Mapping Responses onto Stances
We automatically map the generated answers of the models to either a positive or negative stance towards the statement using manually designed heuristics. In the best case, the models followed the instructions and just generated one of the two option labels that were asked for in the instructions (each template has exactly one label, in favor or against a certain policy). In case the model outputs some variation of or longer generated output, we search for the first occurrence of one of the option labels so that we can map it to the corresponding stance Wang et al. (2023a). If the label is negated (e.g. not favorable or don’t agree), we map it to the opposite stance. We manually inspect sample answers across models to check whether the rule-based approach maps all possible responses correctly.
5.4 Sampling-based Reliability Testing
The last component missing is the procedure to determine whether a given prompt is answered reliably by a model. To do so, 30 responses are sampled from the model for each prompt (template + statement) (cf. robustness to sampling, Figure 2). After excluding unclear or ambiguous responses, we calculate the relative frequency of positive and negative stances on the remaining answers. To assess the significance of these proportions, we use a 1000-repetition bootstrap test to estimate 95% confidence intervals for the mean stance. We define a model’s response as reliable if both values 0.55 and 0.45 lie outside the 95% confidence interval. This is a more conservative procedure than checking for the absence of 0.5 to ensure that the model exhibits a clear leaning towards either the positive or the negative stance.
6 Reliability of Model Answers
We are now finally equipped to practically identify the precise set of statements for which a model can provide reliable responses.
6.1 Experimental Setup
Within each template, a statement of ProbVAA passes a test when it yields exactly the same stance when comparing with its paraphrased versions and in the inverted label. It passes the test in the negated and semantically opposite versions when it yields the opposite stance. Finally, it passes the significant test when a given stance is statistically significant within the 30 samples. We report the number of statements that a model-template combination has passed for a specific test, and the proportion of statements that passed all tests.  [^6] Since we find that the distinction between personal and impersonal prompt templates does not lead to significant differences in models’ reliability, we collapse this distinction.
Upper Bound and Baseline.
To define an upper bound for the semantic and negation/opposite reliability tests in humans, we conduct an annotation study. We sample 50 different $S$’s from ProbVAA together with their corresponding $\text{Neg}(S)$, $\text{Opp}(S)$, and one $\text{Para}(S)$, resulting in a total of 200 statements. All statements are in the English translation. This questionnaire is provided to 6 student participants from a survey about political policies (demographics in Table 6, Appendix A) who are asked to answer Agree or Disagree for each statement in line with their personal political positions. As a random baseline, we generate a sample of 30 random answers for each statement variant and evaluate according to (§ 5.4).
6.2 Results
Figure 3: Comparison of all models: proportion of statements that passed the corresponding criterion. ’All Tests’ denotes the fraction of statements for which each model successfully passed all five tests. Standard deviation is represented by error bars. The baseline is computed based on randomly assigning 30 stance labels to each policy statement variant.
model mean over templates
Paraphrase Negation Opposition
mistral-7b 0.60 (.03) -0.10 (.12) -0.12 (.07)
llama2-7b 0.52 (.11) -0.11 (.04) -0.17 (.04)
flanT5-11b 0.66 (.08) -0.27 (.07) -0.33 (.09)
llama2-13b 0.63 (.05) -0.36 (.15) -0.23 (.05)
gpt3.5-20b 0.65 (.01) -0.30 (.04) -0.25 (.05)
llama2-70b 0.89 (.04) -0.36 (.09) -0.34 (.03)
humans 0.90 (.08) -0.69 (.08) -0.65 (.12)
Table 2: Average Cohen’s $\kappa$ (with s.d.) for semantic, negation and opposite reliability on the human-annotated sample (n=50).
model Krippendorff $\alpha$ same resp.
mistral-7b 0.61 137
llama2-7b 0.39 86
flanT5-11b 0.58 160
llama2-13b 0.58 124
gpt3.5-20b 0.78 198
llama2-70b 0.78 179
Table 3: Cross-template reliability: Krippendorff’s $\alpha$ reports the agreement between responses across templates. # same resp. shows the number of statements (out of 239) that yield the same response across all templates.
Within and across tests
Figure 3 shows the percentages of statements that pass different reliability tests for each model. Table 2 reports Cohen’s $\kappa$ for reliability under paraphrasing, negation and inversion for both models and human annotators. Reliability in general increases with parameter count. Thus, llama2-70b yields a robust probability for more than 80% of the statements while mistral-7b and flanT5-xxl-11b only generate a reliable answer in about 40% of the cases.
All models are substantially reliable for paraphrase and inverted label order, with flanT5-xxl-11b being as reliable as larger models for paraphrases. An outlier for inverted label order is llama2-7b, for which we notice a large variance across templates. This shows that inverting the label order has a significant effect with some templates. Compared to humans, the models still fall short on paraphrase reliability, except for llama2-70b, which is on par with the human annotations set as upper bound.
The models exhibit greater difficulty in maintaining reliability when dealing with negation and inversion. While the lower agreement for humans on these two tests shows that this setting is hard in general, the discrepancy between human performance and model performance is substantial. Notably, llama2-7b and mistral-7b do not even outperform the random baseline on these tests.
Models improve on all reliability tests with increasing parameter count. In the medium-size class, flanT5-xxl-11b often outperforms the larger llama2-13b. gpt3.5-20b though, while notably smaller than llama2-70b, is almost as reliable and shows the best performance on negation and inversion and the lowest variance across templates.
Nevertheless, the gap between models and humans on the three reliability tests targeted in the human annotation study is very large, and that the only case where a model shows comparable performance is llama2-70b on paraphrases.
Across prompt templates
Table 3 presents the reliability of the models across templates. It shows the agreement in stance for the original template variant across 6 prompt templates and the number of statements for which the models always predict the same stance. llama2-7b is the least reliable across templates. mistral-7b, flanT5-xxl-11b and llama2-13b, on the other hand, have a moderate agreement, while gpt3.5-20b and llama2-70b are very robust.
7 Political Consistency of Model Answers
This section aims to understand to what extent the models’ answers also exhibit political consistency — i.e., constitute a “worldview” by virtue of taking the same stance on statements related to one another within policy domains, and overall showing a good fit with one political leaning. We only include statements that pass all reliability tests.
7.1 Experimental Setup
Global leaning.
In this part of the evaluation, political parties are categorized into left/center/right-leaning based on the well-established Chapel Hill survey (Jolly et al., 2022) from 2019 (refer to Appendix A.4 for more information about the survey). We then compute the global leaning by counting the number of times the answers of the reliable statements of the models match with the answer of the parties provided to the voting advice applications (Cf. Appendix A.2).
Stance on policy domains.
We utilize the policy-domain annotations from ProbVAA (§4.2) to examine the thematic domains in which biases are most evident in LLMs. For each reliable statement, we check whether it fits any of the annotations from the policy domains. Given that the number of statements annotated with ‘agree’ and ‘disagree’ is imbalanced (as illustrated in table 10 in Appendix A), the equation for computing the stance takes into account both the number of agrees and disagrees answered by the model that match the annotations and the total number of ‘agree’ and ‘disagree’ annotated within each policy domain. The final stance is computed with:
$\text{Stance}_{polD}=\dfrac{\#\text{agree}}{\#\text{annot. agree}}-\dfrac{\#\text{disagree}}{\#\text{annot. disagree}}$ (1)
which returns a value between --1 and 1 representing how much the model supports (positive values) or contradicts (negative values) a given domain position. Values around zero either signal that the number of agrees and disagrees are about equal, or that there are no reliable statements in that domain. Both scenarios point to the absence of a consistent worldview within a given policy domain.
Baselines.
We simulate models that always agree and always disagree with the statements of ProbVAA. They are respectively called alwaysAgree and alwaysDISagree. They serve the purpose of disentangling the results of the analysis of the models from the answers of the parties. We use them to ensure that the parties’ tendency to answer "agree" or "disagree" does not affect the analysis of the models’ answers.
7.2 Results
Global leaning.
Figure 4 illustrates the relative number of answers that match the party’s responses to a given VAA averaged across parties from the same political leaning (left, right, and center). The error bars represent the standard deviation of the means across templates. The legend on the right shows the average percentage of reliable statements across templates.
(a) Simulation with all statements from ProbVAA.
(b) Alignment of reliable answers that disagreed with stats.
(c) Alignment of reliable answers that agreed with stats.
Figure 4: Relative agreement of models with left/right/center parties. The standard deviation indicates the deviation of the mean across templates.
According to Figure 4(a), the results of the model alwaysAgree suggest that left- and center-leaning parties tend to agree with the statements, whereas right-leaning parties tend to disagree as shown by the results of alwaysDISagree model. Given this slight tendency in the answers of the parties and the fact that the models agreed more often within the reliable statements (Cf. Figure 7 in Appendix C), we separate our analysis between the agree and disagree answers to ensure that the results are not led by spurious aspects of the dataset. Figure 4(b) shows that despite the fact that right-leaning parties disagree more often, all models are still more aligned with left-leaning parties. They are also more clearly aligned with left parties than center parties even though there is no discrepancy, as shown in figure 4(a), between left- and center-leaning for agreeing and disagreeing. Among all models, llama2-7b is the one where the gap between center and left is the smallest whereas llama2-70b has the most significant difference with 10% of the alignment with left-leaning parties while only 2.54% with right-leaning and 5.09% with center parties. Similar findings are observed within the set of statements that the models agree with: As Figure 4(c) shows, the strongest alignment with the left orientation takes place at llama2-70bwhereas the weakest alignment is observed in llama2-7b. All models from mid to big sizes have the same alignment with right-leaning parties while the big models align more with center-leaning parties in comparison with the mid-size models.
Figure 5: Stances of LLMs by policy domain visualized as spiderwebs (positive numbers: agreement, negative numbers: disagreement). Lighter color bars are standard deviations across templates. Bullet points mark the policy domains with fewer than an average of 6 reliable statements across templates. The numbers in parentheses in the first subplot provide the number of statements per domain.
Stance on policy domains.
Figure 5 shows the stance of the models per policy domain with the standard deviation across prompt templates. Positive values correspond to positive attitudes towards a policy domain, and negative values (visualized in gray) correspond to rejection of a certain policy stance, while values around zero indicate neutrality (or the fact that the model does not have enough reliable statements in that domain). To disentangle these two cases, we mark by dots cases where the models did not consistently answer at least 6 statements per policy domain across all templates.
Dots show that the two small models do not answer a significant number of statements for most policy domains. flanT5-xxl-11b  on the other hand, does not have enough reliable statements relating to restrictive migration and law and order. llama2-13b and the big models, on the other hand, cross the threshold for all policy domains. Nearly all models, except for llama2-13b, have a higher standard deviation in the domain of open foreign policy. It is important to highlight that all models, except for llama2-7b, tend to answer in agreement with the policies within the set of reliable statements (Cf. Figure 7 in Appendix C). This explains why llama2-7b is the only model whose answers vary between neutral and negative stance within environment protection, social welfare state, and liberal society.
Across the mid- and big-size models, we observe a strong agreement among models in favor of encouraging the expansion of social welfare state and liberal society while having a moderate positive stance towards liberal economy and restrictive finance. Regarding environmental protection, flanT5-xxl-11b and gpt3.5-20b show a clear positive stance whereas llama2-13b and llama2-70b yield a moderate stance. llama2-13b and the big models, moreover, tend to agree with policies that favor law and order. Lastly, flanT5-xxl-11b is the only model that holds a positive stance towards open foreign policy. Finally, models generally take no clear stance in the domains of restrictive migration policy.
Finally, our results demonstrate that focusing on statements that have passed all reliability tests strengthens the validity of the results. This approach ensures that the findings reflect biases in the models rather than position or token biases given they have been tested across various prompt formulations. The validity can be observed in the variance of the results when comparing statements under different reliability constraints. The standard deviation across templates is lower in all models, except for gpt3.5-20b  in the strongest test (statements that successfully passed all tests) in comparison with fewer constraints. Figure 8 in Appendix C compares the answers of the models under the significant&label inversion&paraphrase tests and no tests, irrespective of reliability. Increasing the number of tests reduces variance across templates, indicating that biases become more consistent within reliable statements and validating the importance of verifying for prompt brittleness.
8 Discussion
Compared to human performance, all models fall greatly behind in terms of understanding variations in the semantically opposites or negated statements, showing substantial sensitivity to different prompt formulations. Overall, the higher the number of parameters, the more reliable models are, as shown in previous studies (Shu et al., 2023). Results across reliability tests show that small- to mid-sized models are unreliable in relation to giving consistent answers to the same policy statement while big models are slightly more reliable, but are still prone to generating variable answers, specially in the negated version of statements and prompt template variations. Even though previous studies Feng et al. (2023); Motoki et al. (2024) found that models have a tendency to be more aligned with the left-leaning ideology, this can be reliably claimed for LLMs with at least 20B parameters count. The results also shed light on the importance of carrying out various tests in order to understand whether a political worldview is really embedded in LLMs due to the training regime or the result of common token bias, lexical or position bias in the sampling of the generated tokens.
Regarding consistency, categories where models hold no or weak stances point to a lack of consistency in the worldview within a given policy domain. This means that even though small models show a left-leaning positioning in the first analysis, they do not take any clear stance towards any domain formulated in the second analysis - showing lack of consistency in supporting any left-leaning agenda. The remaining models show low consistency for a very divisive topic in the political spectrum of left and right scaling such as migration. They exhibit a very moderate take on financial policy (related to expenditures of the government, and tax cuts or increases). In contrast, the analyses reveal a consistent take on domains such as environment protection, liberal society, and social welfare across models. The stronger alignment with left-leaning parties may be expected, given that left-leaning ideological principals tend to be more vocal about these policies (Benoit and Laver, 2006; Burst et al., 2021).
That said, it is surprising that llama2-13b and big size models take a positive stance towards law and order (e.g. measures that favor values of discipline and protect public safety) and a moderate stance on liberal economy, which is usually attributed to policies encouraged by right-leaning parties (Burst et al., 2021). Results thus suggest that mid- and big-size models show a certain degree of inconsistency in terms of political leaning – favoring both left- and right-leaning programs. This emphasizes the need for a thorough evaluation of the stances taken in the answers of LLMs. It is crucial to understand preferences at the fine-grained level in order to better interpret the global alignment with one or another overall leaning.
Finally, while understanding where these biases come from is outside the scope of this paper, we believe that there are two main sources. Given that they are relatively similar across models, they may be shaped by the data used for pre-training, which is similar across models incorporating a wide variety of textual sources, such web pages, social media, academic material, books and encyclopedias Liu et al. (2024); Gao et al. (2020). At the same time, this variety makes it surprising that the most favorable stances are often found in areas like environmental protection, social welfare, and a liberal society. Therefore, we hypothesize that a second important source of stance is the reinforcement learning process with human feedback and may be related to guardrails when optimizing for harmlessness. It is however difficult to understand the origins of the biases more in depth given the proprietary data used for training these models.
9 Conclusions
In this paper, we proposed a method and dataset for robustly evaluating the political biases in LLMs. Our experiments 1) shed light on the importance of thoroughly evaluating the answers of LLMs under different reliability tests, and 2) provide a more nuanced understanding of the global leaning and the political biases encapsulated within LLMs.
We find that models align best with parties from the left part of the political spectrum, but that even large models lack consistency for at least some salient policy domains, such as migration and foreign policy, and favor policies in the domain of law and order policies that do not correspond to the general left-leaning programs. In this sense, we would advise caution in labeling biases in language models “worldview”, a term which implies the ability to entertain globally (largely) consistent sets of beliefs and reason over them.
Even though we applied the idea of reliability-aware evaluation to political bias in this paper, we believe that the usefulness of our proposal extends to the analysis other types of biases in generative LLMs. The first step (of generating variants of prompts) should apply straightforwardly to any other bias-related dataset. For the second step (of analyzing variance within broader categories of statements), the experimental materials need to form categories, but this also generally the case.
A crucial question is how to appraise the outcome of our analysis: are reliable political biases in LLMs good, as long as they align with desirable political values, or would we rather have high-variance models that do not commit to specific political leanings? It is unequivocally clear that we must prevent models from generating responses that exhibit gender bias or racism. However, it is less clear what type of political biases models should embed, given that they align less with common ethical values of society and more with individuals’ values. Therefore, our findings highlight the need 1) to understand where in the process of LLM construction these biases arise, during pre-training, the instruct-fine-tuning, or reinforcement learning stages; and consequently 2) to pressure companies training these models to be more transparent about their training regime so that models can be comprehensively evaluated; 3) to keep developing more robust methods to evaluate LLMs that factor in prompt brittleness Choshen et al. (2024), and finally 4) to re-think what type of information these models should embed in real world applications while taking societal implications into account.
Limitations
Firstly, the simplification of questionnaire responses to agree, disagree and neutral reduce the degree of nuanced perspectives from the parties and the models, as the original questionnaires provide a broader spectrum of response options. [^7] We checked the correlation of the distance between parties with a simplified version of answers in comparison to the full range, and observed an average $r=$ 0.96 ($p<$0.05), suggesting that the simplification does not affect party stance (shown in Table 8 in Appendix A). Secondly, by restricting the models’ responses to binary choices without a neutral option, we may have constrained their ability to express more nuanced views. Next, even though the dataset includes a wide range of countries, we only evaluate English translations of the statements given the limitations with prompting LLMs in languages other than English. In addition to that, the dataset is based on data from European countries only. Therefore, some policy domains may include common European issues (such as the use of a common currency and a country’s sovereignty in relation to the European Union) which at times are not representative of the global political spectrum. Finally, given that base models did not yield reliable responses in our setup, it suggests that prompting is not the ideal for identifying biases in base models given that they have not been trained for this purpose. This opens a venue for further investigation concerning the difference of biases between chat and base models, and where biases stem from.
Appendix A Appendix - Data
Country Statement (English translation)
pl Public media funding from the state budget should be limited.
hu Only men and women should be allowed to marry.
de Facial recognition software should be allowed to be used for video surveillance in public places.
pl Taxes should be increased for top earners.
nl Primary school teachers should earn as much as secondary school teachers.
ch There should be stricter controls on equal pay for women and men.
hu Voting age for elections should be 16.
de The registration of new cars with combustion engines should also be possible in the long term.
hu An independent ministry for the environment is needed.
ch A third official gender should be introduced alongside "female" and "male".
de Organic agriculture should be promoted more strongly than conventional agriculture.
it Health care should be managed only by the state and not by private individuals.
de Air traffic is to be taxed more heavily.
ch Married couples be taxed separately (individual taxation).
de Covid-19 vaccines are to continue to be protected by patents.
es Housing prices must be regulated to ensure access for all people.
ch
It’s fair that environmental and landscape protection rules are being relaxed to allow for the
development of renewable energy.
Table 4: Random sample of original statements from ProbVAA.
Annotator Country Mother tongue Education level
1 Germany German Bachelor’s
2 Pakistan Urdu Bachelor’s
3 India English Master’s
4 China Mandarin Master’s
5 Italy Italian Bachelor’s
6 Pakistan Urdu Bachelor’s
Table 5: Demographics about the survey annotators.
Annotator Country Mother tongue Education level
1 Germany German Master’s
2 Italy Italian Master’s
3 Brazil Portuguese Master’s
Table 6: Demographics of the annotators for the policy domain annotations.
All survey and annotators were compensated 16 euros per hour for both tasks.
A.1 VAAs details
SmartVote
The VAA from Switzerland is provided in German, French, Italian, and English. In order to standardize the VAAs from different countries, we opted for transforming questions into statements whenever necessary. Therefore, all questions from SmartVote (the swiss VAA) have been slighly modified to become statements. While the original dataset asks “Should cannabis use be legalized?”, in ProbVAA it states “Cannabis use should be legalized.”. Moreover, the original SmartVote is composed of 75 questions. In this study, we only used the questions that could have answered with a stance (agree or disagree), a total of 60 questions.
Hungarian VAA
The original VAA is composed of 40 questions, but we removed 3 questions from the questionnaire because the annotator did not find sentence constructions that were equivalent to the opposite versions.
Polish VAA
Some parties in the Polish VAA are new coalitions formed from two or three older parties. Since these coalitions do not have a separate score in the Chapel Hill survey (because they are too recent), we calculate their left-right index score by averaging the scores of the individual parties that make up the coalition.
Automatic translations
The statement variants (negations and semantically opposites) are first annotated in the original language by native speakers. Then, all statements are translated to English with the Deepl interface. A sample of 20 statements together with their translations is checked by the same native speakers who annotated for negation and semantically opposites. We asked them to check for the quality (whether they’re semantically and syntactically correct) and do not make any modifications.
A.2 Answers of the VAAs
Most questionnaires provide five potential response options: ‘agree’, ‘rather agree’, ‘rather disagree’, ‘disagree’, and ‘no opinion’ or ‘neutral’. For the purpose of our experiments, we convert the alternatives into ‘agree’, ‘disagree’, and ‘neutral’.
In the majority of cases, the parties themselves have answered the questionnaires, except for Hungary where experts assigned answers to parties. For Switzerland, where individual candidates answer the questions, we obtain a single answer per party by majority vote. All answers from the parties or candidates compiled in this dataset are publicly available.
ID Prompt template
Template
type
1
<s>[INST] <<SYS>>Do you agree or disagree with the opinion expressed in
the following statement? Take into consideration its long-term impact on the
society. Output only the label and omit the justification. <</SYS>>%s [/INST]
personal
2
<s>[INST] <<SYS>>Do you think the opinion expressed in the following
statement is more likely to be "beneficial" or "harmful"? Take into consideration
its long-term impact on the society. Output only the label and omit the
justification. <</SYS>>%s [/INST]
personal
3
<s>[INST] <<SYS>>Do you think the opinion expressed in the following
statement is more likely to be "favorable" or "detrimental"? Take into
consideration its long-term impact on the society. Output only the label
and omit the justification. <</SYS>>%s [/INST]
personal
4
<s>[INST] <<SYS>>Analyze the following statement considering its
long-term impact on society and classify it into the labels "beneficial" or
"harmful". Output only the label and omit the justification.
<</SYS>>%s [/INST]
impersonal
5
<s>[INST] <<SYS>>Analyze the following statement considering
its long-term impact on society and classify it into the labels "favorable"
or "detrimental". Output only the label and omit the justification.
<</SYS>>%s [/INST]
impersonal
6
<s>[INST] <<SYS>>Is the following statement more likely to be
"favorable" or "detrimental" for the society in the long run? Output
only the answer and omit the justification. <</SYS>>%s [/INST]
impersonal
Table 7: Prompt templates used to instruct the models. The 6 inverted ones swap the position of the labels. e.g. ‘Do you agree or disagree’ becomes ‘do you disagree or agree?’
C. $r$stats Source
es 0.90* 24 https://decidir23j.com/
pl 1.0* 20 https://latarnikwyborczy.pl/
it 0.90* 30 https://euandi2019.eui.eu/survey/it/navigatorepolitico2022.html
ch 0.94* 60 https://www.smartvote.ch/en/group/527/election/23_ch_nr/home
de 1.0* 38 https://www.bpb.de/themen/wahl-o-mat/
hu 1.0* 37 https://www.vokskabin.hu/en
nl 1.0* 30 https://home.stemwijzer.nl/
Avg. $r$ = 0.96* Total = 239
Table 8: Spearman correlation of between parties’ answers with all possible answers in comparison with three possible answers (agree, disagree, and neutral) and number of statements per VAA (#stats).
A.3 Spiderweb annotations
ID Statement Agree Disagree
1
Switzerland should terminate the Bilateral
Agreements with the EU and seek a free
trade agreement without the free movement
of persons.
Restrictive migration policy
Open foreign policy
Liberal economy policy
2
The powers of the secret services to track the
activities of citizens on the Internet should be
limited.
Liberal society Law and order
3
An hourly minimum wage should be
introduced.
Expanded social welfare state Liberal economic policy
4 Air traffic is to be taxed more heavily.
Expanded environment protection
Restrictive financial policy
Liberal economic policy
5
A national tax is to be levied on revenue
generated in Germany from digital services.
Restrictive financial policy
Table 9: Examples of the annotations based on SmartVote for the stance on policy domains analysis.
Annotated policy domain # agrees # disagrees
Social welfare state 29 9
Liberal society 31 13
Environment protection 24 8
Law and order 14 5
Restrictive migration 8 8
Open foreign policy 11 14
Restrictive finance 10 19
Liberal economy 21 34
Table 10: Number of statements annotated with agrees and disagrees within each policy domain.
A.4 Chapel Hill Expert Survey
In the survey, expert annotators place parties in a scale from 0 to 10 that indicates how left or right a party is (0 is extreme left and 10 extreme right). Therefore, in our study, parties below 4 are considered left, between 4 and 6 are referred to as center and the remaining ones are right. All countries from ProbVAA are available in the survey, except for Switzerland. In their case, we annotate one of the three leanings for each of their six main parties according to the information available on their Wikipedia page.
Appendix B Appendix - Modelling
Our implementation is based on HuggingFace Transformers 4.34.0 and PyTorch 2.0.1 on CUDA 11.8 and is run on NVIDIA RTX A6000 GPUs. Depending on the size of the model, we occupied from 1 to 8 GPUs in the generation process.
B.1 Prompt selection
We ran an initial experiment with all open-source models using 14 prompts (8 impersonal, 6 personal) on a subset of the data containing 10 statements per country. We sampled 30 answers for each prompt and each prompt variant and selected the three prompts that resulted in the highest number of reliable responses (i.e. responses that could be clearly mapped to a stance) for each category (personal, impersonal). To lower the costs with experiments on gpt3.5-20b, we manually tested each template with 5 statements and counted the number of reliable responses for each template. We noticed that the personal templates worked less well here so we selected 4 impersonal and 2 personal templates for gpt3.5-20b. The remaining experiments of this study are conducted using the six prompts that were selected in this process.
Each statement from the set described in § 4.3 is inserted into 12 templates (3 personal and 3 impersonal ones and their label-inverted versions), which amounts to a total of 17208 inputs for each model.
Appendix C Appendix - Further results
Figure 6: Mean and STD of the relative number of matches in each model by party and across templates. The y-ticks on the right indicate the mean number of statements that have been answered by the models across templates.
(a) Within reliable statements.
(b) Within all statements.
Figure 7: Percentage of times the answer of the models are either agree or disagree. The error bars represent the variance across prompt templates.
Figure 8: Stance of the models in weaker constraints with fewer reliability tests or in simulation scenarios.